# A Stochastic Genetic Interacting Particle Method for Reaction-Diffusion-Advection Equations

### v1

@@ 5 Conclusion
We have presented the SGIP method, a novel stochastic genetic interacting particle algorithm for solving reaction-diffusion-advection equations. The method’s effectiveness stems from its integrated design that combines three key components: operator splitting for process decoupling, stochastic particle dynamics for transport, and adaptive resampling for long-term stability.
The main theoretical contribution is a rigorous convergence analysis, where we prove that under appropriate assumptions, the method achieves consistent accuracy across various dimensionalities and problem configurations. This provides a solid mathematical foundation for the proposed methodology. Numerical experiments demonstrate SGIP’s capability in handling diverse challenges: from various reaction types (FKPP, cubic, Arrhenius) to complex flow configurations (shear, cellular, ABC flows) across 1D, 2D, and 3D domains. The Lagrangian formulation naturally handles sharp gradients and complex geometries without requiring specialized discretization techniques. Particularly in higher dimensions, the method offers computational advantages over traditional Eulerian approaches by avoiding the curse of dimensionality while maintaining accuracy with relatively large time steps.
The adaptive resampling strategy proves crucial for maintaining numerical stability and computational efficiency over long-time integration. Through intelligent particle redistribution guided by Eulerian density reconstruction, SGIP maintains an optimal balance between statistical accuracy and computational efficiency. Future work will focus on extending SGIP to multi-species systems, and implementing the fast adaptive algorithm discussed in Remark 1 to concentrate computational resources in active transition regions for enhanced computational efficiency in higher dimensions.
6 Acknowledgements
ZW was partly supported by NTU SUG-023162-00001, MOE AcRF Tier 1 Grant RG17/24. JX was partly supported by NSF grants DMS-2219904 and DMS-2309520, a Qualcomm Gift Award, and the Swedish Research Council grant no. 2021-06594 during his residence at the Institut Mittag-Leffler in Djursholm, Sweden, and the support during his stay at the E. Schrödinger Institute, Vienna, Austria, in the Fall of 2025. ZZ was supported by the National Natural Science Foundation of China (Projects 92470103 and 12171406), the Hong Kong RGC grant (Projects 17304324 and 17300325), the Seed Funding Programme for Basic Research (HKU), the Outstanding Young Researcher Award of HKU (2020–21), and the Seed Funding for Strategic Interdisciplinary Research Scheme 2021/22 (HKU).

### v2

@@ 5 Conclusion
We have presented the SGIP method, a novel stochastic genetic interacting particle algorithm for solving reaction-diffusion-advection equations. The method’s effectiveness stems from an integrated design that combines operator splitting for process decoupling, stochastic particles for transport, and adaptive resampling for long-term stability.
The main theoretical contribution is a complete convergence analysis, proving consistent accuracy of the method across various dimensionalities. Numerical experiments demonstrate SGIP’s capability in handling diverse challenges, including different reaction types (FKPP, cubic, Arrhenius) and complex flows (shear, cellular, ABC) in 1D to 3D domains. The Lagrangian formulation naturally resolves sharp gradients and complex flow geometries without resorting to specialized discretizations. Notably, in higher dimensions, the method offers computational advantages by mitigating the curse of dimensionality and maintaining accuracy with relatively large time steps.
The adaptive resampling strategy is crucial for sustained stability and efficiency. By intelligently redistributing particles based on Eulerian density reconstruction, it balances statistical accuracy with computational cost. Future work will extend SGIP to multi-species systems and implement the fast adaptive algorithm discussed in Remark 2.1 to concentrate computational resources in active transition regions, thereby enhancing efficiency for higher-dimensional problems.
Acknowledgements
ZW was partly supported by NTU SUG-023162-00001, MOE AcRF Tier 1 Grant RG17/24. JX was partly supported by NSF grants DMS-2219904 and DMS-2309520, a Qualcomm Gift Award, and the Swedish Research Council grant no. 2021-06594 during his residence at the Institut Mittag-Leffler in Djursholm, Sweden, and the support during his stay at the E. Schrödinger Institute, Vienna, Austria, in the Fall of 2025. ZZ was partly supported by the National Natural Science Foundation of China (Projects 12171406 and 92470103), the Hong Kong RGC grant (Projects 17304324 and 17300325), the Seed Funding Programme for Basic Research (HKU), and the Hong Kong RGC Research Fellow Scheme 2025. The computations were performed using the HKU ITS research computing facilities.